find the probability that a person tossing 8 coins will get either all heads or all tails for the first time on the seventh toss?
1
Expert's answer
2020-02-10T10:45:08-0500

Probability of getting 8 heads (tails) in one toss: "(\frac{1}{2})^8=\frac{1}{256}"

Probability of getting 8 heads or tails in one toss: "\frac{1}{256}*2=\frac{1}{128}"

Probability of getting 8 heads or tails for the first time on seventh toss: "(1-\frac{1}{128})^6*\frac{1}{128}=\frac{127^6}{128^7}\approx 2.26*10^{-13}"